# Conceptualization of this study, Methodology, Software

### v1

@@ 2.3.3 Myenteric Plexus
Similar to the training procedure for segmenting the muscularis propria, a 5-fold cross-validation strategy was also applied throughout training to minimize the risk of the pre-trained ViT model overfitting to the dataset. This was achieved by utilizing five different folds, each consisting of 6 WSIs. Therefore, each WSI was used for both training and testing, thereby allowing the model to make full use of each WSI. Identical data augmentation techniques were applied when preparing the tiles for training purposes. This also included scaling, a 90 random rotation, and horizontal and vertical flipping of approximately 180, each with a 50% chance of occurrence. To mitigate any potential effects that rotation operation could have had on performance, the border edges of the binary mask were trimmed to preserve alignment between the original tiles and ground truth tiles, specifically in regions of interest. The ViT-based model also utilized an AdamW optimizer. Unlike the segmentation of the muscularis propria, segmenting the myenteric plexus utilized a learning rate of 2e-4 and a weight decay of 1e-3. Similarly, a cosine learning rate scheduler was also used with five warmup epochs followed by a full training session of 50 epochs.
@@ 2.3.4 Ganglion Cells
Similar to the previous segmentation steps, a 5-fold cross-validation scheme was employed to maximize data use while preventing overfitting. Five folds were defined, each with six WSIs. This procedure was repeated for all five folds, such that every WSI was used as a test sample exactly once and was never present in the training set of the same fold.
@@ 2.5.1 Muscularis Propria
For muscularis segmentation, performance was assessed using the Dice coefficient, PIR, precision, and recall. The Dice coefficient measured the overlap between the predicted and ground-truth muscularis regions. At the same time, PIR quantified how effectively the segmented muscularis retained the plexus regions, which is necessary for downstream layers. Precision and recall quantified pixel-level accuracy. A threshold sweep identified 0.01 as the optimal threshold, producing the highest Dice coefficient and PIR (Figure 5).
@@ 2.5.2 Myenteric Plexus
For myenteric plexus segmentation, precision, recall, and GIR were used. Precision measured how many predicted plexus pixels were correct, while recall measured how completely the plexus regions were captured. GIR quantified how many ganglion-containing plexus regions were retained, which directly affects the success of ganglion detection. The optimal threshold identified for plexus segmentation was 0.25.
@@ 2.5.3 Ganglion Cells
For ganglion cell detection, precision and recall were used to evaluate object-level accuracy. Precision represented the proportion of predicted ganglion cells that were correctly identified, while recall measured how completely the model captured the ganglion cells present in the ground truth. A threshold sweep determined that 0.175 provided the best balance between precision and recall for ganglion.
2.6 Baseline Model
@@ 2.6.1 Muscularis Propria
The baseline model’s methodology to successfully segment the muscularis propria consisted of utilizing Calretinin-stained WSI of the colon cross-sections and deploying a colour-based k-means clustering algorithm to segment the colon section [14]. Additionally, a CNN-based algorithm that was pretrained on ImageNet-1k was also utilized to segment the muscularis propria layer [13, 14], using tiles with dimensions of 256256 pixels. Postprocessing techniques were also applied, which included morphological smoothing [19]. The segmentation performance was evaluated utilizing the Dice coefficient, PIR, in addition to precision and recall.
@@ 2.6.2 Myenteric Plexus
The segmentation of the myenteric plexus involved receiving the binary masks, corresponding to the model’s prediction output, and utilizing these images for the segmentation process. Unlike the muscularis propria segmentation, the myenteric plexus was segmented first, using a colour-based thresholding to differentiate between brown-stained areas [14]. This was then followed by applying a pixel threshold to accept and reject pixels that could correspond to being a part of the myenteric plexus layer. Furthermore, morphological filtering [19] was also applied to filter out pixels that potentially correspond to true or false myenteric plexus regions for the final stage of segmentation. To evaluate the segmentation accuracy of the myenteric plexus layer, the GIR was also utilized, alongside precision and recall.
@@ 3 Results
The segmentation of the muscularis propria region is of significant importance, as it shapes the anatomical boundaries for the later identification of ganglion cells located within the myenteric plexus. Therefore, isolating the muscularis layer allows segmentation of the plexus region, where ganglion cells are typically found. Furthermore, this narrows the region of interest from a much broader WSI to only tissue regions clinically more relevant to diagnosing HD, as evidenced by a high Dice coefficient. Moreover, confidence thresholds, which determine the model’s ability to include a tile as part of the muscularis propria, were experimented with. A confidence threshold of 0.4 (Figure 5) yields model results that are similar to those of the ground truth. However, a lower confidence threshold of 0.01 achieves a superior PIR with a lower Dice coefficient, giving further certainty that all tiles include a plexus region, which serves the ultimate goal of detecting ganglion cells better, as shown in Figure 5.
@@
The results, as shown in Tables 1, demonstrate a high Dice coefficient of 89.9%, indicating a more accurate segmentation performance of the muscularis layer compared to the ground truth manual segmentation. This not only shows a high overall accuracy but also superior performance compared to the baseline CNN and K-means models. Additionally, the model achieved a precision score of 82.4%, indicating that the muscularis region corresponds to muscularis tissue, mislabeling only a few non-muscularis tissues. The model also achieved a recall score of 99.7%, demonstrating its superior ability to identify all positive tissue pixels. The model also achieved a PIR of 100%, which was deemed highly important, as all ganglion cells that needed to be identified later resided within the myenteric plexus. That said, all model-predicted outputs contained plexus regions, which play an instrumental role in the subsequent segmentation stage. A qualitative comparison of predicted muscularis segmentation with plexus overlays is shown in Figure 6.
| Model | Precision | Recall | GIR |
| Colour-based thresholding [11, 14] | 73.8 | 85.9 | 99.2 |
| ViT-B/16 (proposed) | 84.2 | 94.8 | 99.7 |
@@
The following segmentation stage included segmenting the myenteric plexus section of the colon. This stage is considered the most crucial segmentation phase before ganglion identification, as ganglion cells are located within the myenteric plexus layer. The ViT-B/16 proposed model achieves high segmentation accuracy in the plexus layer, which prepares for optimal ganglia identification in the downstream task. A confidence threshold of 0.25 was used to select the predicted pixels in the binary output prediction map. Although the chosen confidence threshold was flexible, it yielded a higher recall at the expense of precision, prioritizing a higher GIR score. This threshold yielded the optimal GIR, as well as precision and recall. A representative plexus prediction example with ganglion cell annotations is shown in Figure 7.
| No. | WSI Name | TP | FP | FN | Precision | Recall | GIR |
| 1 | S14-580 | 158 | 55 | 17 | 0.742 | 0.903 | 0.978 |
| 2 | S00-1910 | 79 | 5 | 1 | 0.940 | 0.988 | 1.0 |
| 3 | S02-410 | 101 | 8 | 2 | 0.927 | 0.981 | 1.0 |
| 4 | S02-484 | 141 | 24 | 0 | 0.855 | 1.000 | 1.0 |
| 5 | S03-2391 | 96 | 7 | 9 | 0.932 | 0.914 | 1.0 |
| 6 | S01-18 | 182 | 32 | 9 | 0.850 | 0.953 | 1.0 |
| 7 | S03-3178 D2 | 63 | 8 | 1 | 0.887 | 0.984 | 1.0 |
| 8 | S03-3178 D3 | 47 | 16 | 0 | 0.746 | 1.000 | 1.0 |
| 9 | S03-3178 D4 | 59 | 6 | 6 | 0.908 | 0.908 | 1.0 |
| 10 | S04-52 | 82 | 16 | 2 | 0.837 | 0.976 | 1.0 |
| 11 | S04-910 | 156 | 33 | 8 | 0.825 | 0.951 | 1.0 |
| 12 | S07-1808 | 124 | 30 | 1 | 0.805 | 0.992 | 1.0 |
| 13 | S08-2215 | 161 | 26 | 10 | 0.861 | 0.942 | 1.0 |
| 14 | S09-2723 | 43 | 21 | 7 | 0.672 | 0.860 | 1.0 |
| 15 | S04-1840 | 118 | 13 | 4 | 0.901 | 0.967 | 0.976 |
| 16 | S07-1465 | 162 | 37 | 17 | 0.814 | 0.905 | 1.0 |
| 17 | S14-1715 | 62 | 12 | 2 | 0.838 | 0.969 | 1.0 |
| 18 | S09-2909 | 71 | 20 | 0 | 0.780 | 1.000 | 1.0 |
| 19 | S14-3414 | 108 | 8 | 5 | 0.931 | 0.956 | 1.0 |
| 20 | S14-2038 | 164 | 33 | 2 | 0.832 | 0.988 | 1.0 |
| 21 | S15-1442 | 127 | 23 | 0 | 0.847 | 1.000 | 1.0 |
| 22 | S15-1518 | 111 | 9 | 4 | 0.925 | 0.965 | 1.0 |
| 23 | S16-567 | 118 | 34 | 4 | 0.776 | 0.967 | 1.0 |
| 24 | S16-1197 B1 | 51 | 16 | 0 | 0.761 | 1.000 | 1.0 |
| 25 | S11-1760 | 223 | 14 | 17 | 0.941 | 0.929 | 1.0 |
| 26 | S16-1467 | 84 | 10 | 3 | 0.894 | 0.966 | 0.950 |
| 27 | S16-1197 B3 | 70 | 8 | 8 | 0.897 | 0.897 | 1.0 |
| 28 | S16-1197 B2 | 100 | 35 | 1 | 0.741 | 0.990 | 1.0 |
| 29 | S97-2054 | 90 | 6 | 53 | 0.938 | 0.629 | 1.0 |
| 30 | S16-1415 | 36 | 18 | 2 | 0.667 | 0.947 | 1.0 |
@@
The results, as shown in Tables 2, demonstrate a high recall score of 94.8%, indicating the model’s ability to detect almost all positive samples, corresponding to regions that contain myenteric plexus. Furthermore, a high precision score of 84.2% demonstrates the model’s capability of detecting regions that not only accurately pertain to the myenteric plexus but also limit false positives by being confined to regions enclosed by the muscularis. Finally, the ViT model achieved a high GIR score of 99.7% when segmenting the myenteric plexus. Given the myenteric plexus’ thin and long anatomical structure, and its thin, elongated, and discontinuous nature, the model’s segmentation performance, with high GIR ( Figure 7), precision, and recall, allows for higher chances of success in later ganglion cell identification. A detailed breakdown of per-slide plexus segmentation outcomes is provided in Table 3.
@@
Once the myenteric plexus was segmented, the next stage involved segmenting and identifying ganglion cells to further enhance clinical relevance to HD. This segmentation stage involved transitioning from region-level segmentation to cell-level (object) prediction. The threshold utilized for segmenting and detecting the ganglion cells is 0.175. An example visualization of predicted high- and low-certainty ganglion cells across a whole slide is shown in Figure 8.
@@
The ViT-B/16 proposed model achieved a high recall average score of 78.6% (Table 4) by utilizing the combined low-certainty high-certainty ganglion cells, demonstrating strong capabilities in detecting the majority of ganglion cells within the myenteric plexus region. Furthermore, using a more conservative threshold for higher certainty yields a precision of 62.1%, indicating the model’s ability to identify clear, morphologically distinct ganglion cells. Using the same high-certainty threshold, the model achieved a high recall of 89.1%, thereby identifying the overwhelming majority of ganglion cells with a typical ganglion-cell morphology. Per-slide high-certainty ganglion identification results are summarized in Table 5. Combined high and low certainty object-level results for each WSI are listed in Table 6. Overall, the DL model used for segmenting and identifying ganglion cells outperformed the baseline model in ganglion cell detection.
| No. | WSI Name | TP | FP | FN | Precision | Recall |
| 1 | S14-580 | 83 | 34 | 9 | 0.709 | 0.902 |
| 2 | S00-1910 | 4 | 2 | 0 | 0.667 | 1.0 |
| 3 | S02-410 | 23 | 17 | 0 | 0.575 | 1.0 |
| 4 | S02-484 | 32 | 46 | 0 | 0.410 | 1.0 |
| 5 | S03-2391 | 13 | 9 | 4 | 0.591 | 0.765 |
| 6 | S03-3178 D2 | 5 | 2 | 0 | 0.714 | 1.0 |
| 7 | S03-3178 D3 | 3 | 9 | 0 | 0.250 | 1.0 |
| 8 | S03-3178 D4 | 1 | 7 | 0 | 0.125 | 1.0 |
| 9 | S04-52 | 8 | 11 | 8 | 0.421 | 0.50 |
| 10 | S04-910 | 30 | 7 | 0 | 0.811 | 1.0 |
| 11 | S07-1808 | 82 | 33 | 0 | 0.713 | 1.0 |
| 12 | S08-2215 | 50 | 18 | 1 | 0.735 | 0.980 |
| 13 | S09-2723 | 22 | 24 | 2 | 0.478 | 0.917 |
| 14 | S04-1840 | 41 | 4 | 0 | 0.911 | 1.0 |
| 15 | S07-1465 | 32 | 74 | 0 | 0.302 | 1.0 |
| 16 | S14-1715 | 4 | 1 | 6 | 0.800 | 0.40 |
| 17 | S09-2909 | 43 | 41 | 1 | 0.512 | 0.977 |
| 18 | S14-3414 | 50 | 17 | 3 | 0.746 | 0.943 |
| 19 | S14-2038 | 70 | 30 | 2 | 0.700 | 0.972 |
| 20 | S15-1442 | 82 | 35 | 1 | 0.701 | 0.988 |
| 21 | S15-1518 | 63 | 20 | 16 | 0.759 | 0.797 |
| 22 | S16-567 | 117 | 23 | 29 | 0.836 | 0.801 |
| 23 | S16-1197 B1 | 51 | 14 | 0 | 0.785 | 1.0 |
| 24 | S11-1760 | 24 | 19 | 2 | 0.558 | 0.923 |
| 25 | S16-1467 | 19 | 4 | 1 | 0.826 | 0.950 |
| 26 | S16-1197 B3 | 35 | 8 | 3 | 0.814 | 0.921 |
| 27 | S16-1197 B2 | 32 | 5 | 0 | 0.865 | 1.0 |
| 28 | S97-2054 | 0 | 1 | 4 | 0.0 | 0.0 |
| 29 | S16-1415 | 22 | 2 | 0 | 0.917 | 1.0 |
| 30 | S01-18 | 10 | 15 | 0 | 0.400 | 1.0 |
| No. | WSI Name | TP | FP | FN | Precision | Recall |
| 1 | S14-580 | 86 | 31 | 14 | 0.735 | 0.860 |
| 2 | S00-1910 | 5 | 1 | 1 | 0.833 | 0.833 |
| 3 | S02-410 | 26 | 15 | 2 | 0.634 | 0.929 |
| 4 | S02-484 | 40 | 37 | 2 | 0.519 | 0.952 |
| 5 | S03-2391 | 13 | 9 | 6 | 0.591 | 0.684 |
| 6 | S03-3178 D2 | 5 | 2 | 2 | 0.714 | 0.714 |
| 7 | S03-3178 D3 | 4 | 8 | 0 | 0.333 | 1.0 |
| 8 | S03-3178 D4 | 4 | 4 | 2 | 0.50 | 0.667 |
| 9 | S04-52 | 12 | 7 | 13 | 0.632 | 0.480 |
| 10 | S04-910 | 34 | 3 | 3 | 0.919 | 0.919 |
| 11 | S07-1808 | 90 | 25 | 8 | 0.783 | 0.918 |
| 12 | S08-2215 | 56 | 13 | 3 | 0.812 | 0.949 |
| 13 | S09-2723 | 25 | 22 | 7 | 0.532 | 0.781 |
| 14 | S04-1840 | 44 | 1 | 13 | 0.978 | 0.772 |
| 15 | S07-1465 | 37 | 69 | 1 | 0.349 | 0.974 |
| 16 | S14-1715 | 4 | 1 | 10 | 0.80 | 0.286 |
| 17 | S09-2909 | 46 | 40 | 12 | 0.535 | 0.793 |
| 18 | S14-3414 | 51 | 16 | 3 | 0.761 | 0.944 |
| 19 | S14-2038 | 74 | 27 | 22 | 0.733 | 0.771 |
| 20 | S15-1442 | 86 | 32 | 6 | 0.729 | 0.935 |
| 21 | S15-1518 | 72 | 11 | 25 | 0.867 | 0.742 |
| 22 | S16-567 | 120 | 20 | 37 | 0.857 | 0.764 |
| 23 | S16-1197 B1 | 55 | 10 | 14 | 0.846 | 0.797 |
| 24 | S11-1760 | 27 | 16 | 15 | 0.628 | 0.643 |
| 25 | S16-1467 | 19 | 4 | 3 | 0.826 | 0.864 |
| 26 | S16-1197 B3 | 39 | 4 | 8 | 0.907 | 0.830 |
| 27 | S16-1197 B2 | 32 | 5 | 5 | 0.865 | 0.865 |
| 28 | S97-2054 | 0 | 1 | 6 | 0.0 | 0.0 |
| 29 | S16-1415 | 22 | 2 | 2 | 0.917 | 0.917 |
| 30 | S01-18 | 14 | 11 | 0 | 0.560 | 1.0 |
@@ 4.1 Interpretation of Results
The muscularis propria is a smooth muscle layer embedded within the walls of organs such as the GI tract and bladder, composed of an inner circular layer and an outer longitudinal layer, giving it a more uniform structure [26]. The ViT has an inherent self-attention mechanism, enabling it to attend to other image patches and capture long-range dependencies. For instance, it could relate to the appearance of the muscularis propria and its location relative to different regions of the image that correspond to the lamina propria or the submucosa. The segmentation performance of the ViT model yielded a Dice coefficient of 89.9%, indicating strong agreement between the predicted and ground truth maps. These results demonstrate that global-context information, combined with local feature representations, enables ViTs to achieve high segmentation accuracy, as evidenced by the Dice coefficients obtained. Additionally, segmentation performance on this layer yields a PIR of 100%, ensuring that the plexus segmentation is applied within the anatomical region of interest for later accurate segmentation of ganglion cells.
@@
The myenteric plexus region is embedded within the inner layer of the muscularis propria and consists of nerve fibres responsible for several autonomous digestive functions. Unlike the muscularis propria, the myenteric plexus is a smaller, less structured region and thus appears more fragmented across WSIs [26, 27], making segmentation more difficult. The myenteric plexus is of great importance because it is the anatomical site where ganglion cells are present [27], which are needed for future segmentation and identification. The segmentation performance yielded recall and precision scores of 94.8% and 84.2%, respectively. Indicating the model’s robust capability of segmenting most of the pixels that correspond to the plexus region. Additionally, among the tiles segmented, fewer false plexus segmentations (false positives) are present. Furthermore, the model produced a GIR of 99.7%, indicating that it not only retained accurate segmentations but also identified relevant areas containing ganglion cells. Because the myenteric plexus is embedded within the muscularis propria, producing textural variability [26], this allows ViTs to leverage their inherent self-attention mechanisms, achieving superior results. For instance, the attention layers embedded within the ViT enable the model to distinguish between the thick, continuous, and homogeneous fibres of the muscularis layer and the thin, irregular fibres of the plexus, even when the muscularis propria and the myenteric plexus share the exact colour tones [28]. The ViT computes attention scores between patch pairs, enabling it to identify which patches are similar or different based on the embeddings it produces. This allows the model to distinguish between slight variations in texture. Furthermore, given the fragmented myenteric plexus throughout the WSI, it is discontinuous across several patches, providing ViTs with an opportunity to leverage their inherent understanding of global contextual relationships. Unlike conventional CNNs, which primarily rely on local receptive fields, this may prevent recognizing that spatially separated patches belong to the same structure, leading to fragmented predictions, especially when there are mild structural differences between patches. However, since ViTs operate globally, the model can label patches with slightly different structures as belonging to the plexus region, even when their embeddings are similar. That being said, high segmentation accuracy and high GIR of the myenteric plexus are critical to optimizing ganglion cell segmentation and identification.

### v2

@@ 2.3.3 Myenteric Plexus
Similar to the training procedure for segmenting the muscularis propria, a 5-fold cross-validation strategy was also applied throughout training (24 WSIs, 24,000 tiles) to minimize the risk of the pre-trained ViT model overfitting to the dataset. Identical data augmentation techniques were applied when preparing the tiles for training purposes. This also included scaling, a 90 random rotation, and horizontal and vertical flipping of approximately 180, each with a 50% chance of occurrence. To mitigate any potential effects that rotation operation could have had on performance, the border edges of the binary mask were trimmed to preserve alignment between the original tiles and ground truth tiles, specifically in regions of interest. The ViT-based model for segmentation of the myenteric plexus also utilized an AdamW optimizer. Unlike the segmentation of the muscularis propria, segmenting the myenteric plexus utilized a learning rate of 2e-4 and a weight decay of 1e-3. Similarly, a cosine learning rate scheduler was also used with five warmup epochs followed by a full training session of 50 epochs.
@@ 2.3.4 Ganglion Cells
Similar to the muscularis propria and plexus segmentation tasks, ganglion cell detection was performed using a 5-fold cross-validation strategy at the WSI level. In each fold, 1000 tiles were randomly extracted from each training WSI and used for model optimization, while tiles from the held-out WSIs were reserved for validation and WSI-level evaluation. Final performance was assessed by re-stitching overlapping tile predictions to form full-resolution WSI segmentation maps.
@@ 2.5.1 Muscularis Propria
For muscularis segmentation, performance was assessed using the Dice coefficient, PIR, precision, and recall. The Dice coefficient measured the overlap between the predicted and ground-truth muscularis regions. At the same time, PIR quantified how effectively the segmented muscularis retained the plexus regions, which is necessary for downstream layers. Precision and recall quantified pixel-level accuracy.
@@ 2.5.2 Myenteric Plexus
For myenteric plexus segmentation, precision, recall, and GIR were used. Precision measured how many predicted plexus pixels were correct, while recall measured how completely the plexus regions were captured. GIR quantified how many ganglion-containing plexus regions were retained, which directly affects the success of ganglion detection.
@@ 2.5.3 Ganglion Cells
Object-level accuracy was calculated for ganglion cell detection using precision and recall, defined in Equations (1) and (2). In this case, a TP is a predicted ganglion object that has an overlap with a ground truth ganglion annotation. A FP is a predicted ganglion object that has no overlap with any ground truth ganglion, and a FN is a ground truth ganglion that was not detected by the model. Precision, then, can be interpreted as how many of the predicted ganglion objects are correct, while recall indicates how completely the model is capturing the annotated ganglion cells.
2.6 Baseline Method
To contextualize the performance of the proposed ViT-based pipeline, results were compared against previously published methods that evaluated the same WSI dataset [11, 13, 14]. These baseline approaches are summarized below.
@@ 2.6.1 Muscularis Propria
The baseline model’s methodology to successfully segment the muscularis propria consisted of utilizing Calretinin-stained WSI of the colon cross-sections and deploying a colour-based k-means clustering algorithm to segment the colon section [14]. Additionally, a CNN-based algorithm that was pretrained on ImageNet-1k was also utilized to segment the muscularis propria layer [13, 14], using tiles with dimensions of 256256 pixels. Postprocessing techniques were also applied, which included morphological smoothing [19].
@@ 2.6.2 Myenteric Plexus
The segmentation of the myenteric plexus involved receiving the binary masks, corresponding to the model’s prediction output, and utilizing these images for the segmentation process. Unlike the muscularis propria segmentation, the myenteric plexus was segmented first, using a colour-based thresholding to differentiate between brown-stained areas [14]. This was then followed by applying a pixel threshold to accept and reject pixels that could correspond to being a part of the myenteric plexus layer. Furthermore, morphological filtering [19] was also applied to filter out pixels that potentially correspond to true or false myenteric plexus regions for the final stage of segmentation.
@@ 3.1 Muscularis Propria
The segmentation of the muscularis propria region is of significant importance, as it shapes the anatomical boundaries for the later identification of ganglion cells located within the myenteric plexus. Therefore, isolating the muscularis layer allows segmentation of the plexus region, where ganglion cells are typically found. Furthermore, this narrows the region of interest from a much broader WSI to only tissue regions clinically more relevant to diagnosing HD, as evidenced by a high Dice coefficient. Moreover, confidence thresholds, which determine the model’s ability to include a tile as part of the muscularis propria, were experimented with. In terms of determining the most optimal threshold to accept pixels corresponding to a muscularis propria, a threshold sweep of 0.01-0.99 was applied, and the threshold which produced the highest PIR and Dice coefficient was utilized. The effect of varying the prediction threshold on both Dice and PIR is summarized in Fig. 5. A confidence threshold of 0.4 yields model results that are similar to those of the ground truth. However, a lower confidence threshold of 0.01 achieves a superior PIR with a lower Dice coefficient (chosen as the optimal value), giving further certainty that all tiles include a plexus region, which serves the ultimate goal of detecting ganglion cells better, as shown in Fig. 5.
@@
The results, as shown in Table 1, demonstrate a high Dice coefficient of 89.9%, indicating a more accurate segmentation performance of the muscularis layer compared to the ground truth manual segmentation. This not only shows a high overall accuracy but also superior performance compared to the baseline CNN and K-means models. Additionally, the model achieved a precision score of 82.4%, indicating that the muscularis region corresponds to muscularis tissue, mislabeling only a few non-muscularis tissues. The model also achieved a recall score of 99.7%, demonstrating its superior ability to identify all positive tissue pixels. The model also achieved a PIR of 100%. That said, all model-predicted outputs contained plexus regions, which play an instrumental role in the subsequent segmentation stage. A qualitative comparison of predicted muscularis segmentation with plexus overlays is shown in Fig. 8.
3.2 Myenteric Plexus
@@
The following segmentation stage included segmenting the myenteric plexus section of the colon. This stage is considered the most crucial segmentation phase before ganglion identification, as ganglion cells are located within the myenteric plexus layer. The ViT-B/16 proposed model achieves high segmentation accuracy in the plexus layer, which prepares for optimal ganglia identification in the downstream task. A confidence threshold of 0.25 was used to select the predicted pixels in the binary output prediction map. Although the chosen confidence threshold was flexible, it yielded a higher recall at the expense of precision, prioritizing a higher GIR score (Fig. 6). This threshold yielded the optimal GIR, as well as precision and recall. A representative plexus prediction example with ganglion cell annotations is shown in Fig. 9.
@@
The results, as shown in Table 2, demonstrate a high recall score of 94.8%, indicating the model’s ability to detect almost all positive samples, corresponding to regions that contain myenteric plexus. Furthermore, a high precision score of 84.2% demonstrates the model’s capability of detecting regions that not only accurately pertain to the myenteric plexus but also limit false positives by being confined to regions enclosed by the muscularis. Finally, the ViT model achieved a high GIR score of 99.7% when segmenting the myenteric plexus. Given the myenteric plexus’ thin and long anatomical structure, and its thin, elongated, and discontinuous nature, the model’s segmentation performance, with high GIR ( Fig. 9), precision, and recall, allows for higher chances of success in later ganglion cell identification.
3.3 Ganglion Cells
@@
Once the myenteric plexus was segmented, the next stage involved segmenting and identifying ganglion cells to further enhance clinical relevance to HD. This segmentation stage involved transitioning from region-level segmentation to cell-level (object) prediction. The threshold utilized for segmenting and detecting the ganglion cells is 0.175 (Fig. 7). An example visualization of predicted high- and low-certainty ganglion cells across a whole slide is shown in Fig. 10.
@@
The ViT-B/16 proposed model achieved a high recall average score of 78.6% (Table 3) by utilizing the combined low-certainty high-certainty ganglion cells, demonstrating strong capabilities in detecting the majority of ganglion cells within the myenteric plexus region. Furthermore, using a more conservative threshold for higher certainty yields a precision of 62.1%, indicating the model’s ability to identify clear, morphologically distinct ganglion cells. Using the same high-certainty threshold, the model achieved a high recall of 89.1%, thereby identifying the overwhelming majority of ganglion cells with a typical ganglion-cell morphology. Overall, the DL model used for segmenting and identifying ganglion cells outperformed the baseline model in ganglion cell detection.
@@ 4.1 Interpretation of Results
The muscularis propria is a smooth muscle layer embedded within the walls of organs such as the GI tract and bladder, composed of an inner circular layer and an outer longitudinal layer, giving it a more uniform structure [26]. The ViT has an inherent self-attention mechanism, enabling it to attend to other image patches and capture long-range dependencies. For instance, it can relate the appearance of the muscularis propria to its spatial context relative to adjacent layers, such as the lamina propria or submucosa. The segmentation performance of the ViT model yielded a Dice coefficient of 89.9%, indicating strong agreement between the predicted and ground truth maps. Notably, the model slightly oversegments the muscularis, reflected in higher recall and lower precision. This behaviour increases the number of false positive muscularis pixels, but it preserves a perfect PIR, ensuring that all plexus regions remain fully contained within the predicted muscularis mask. While this enlarges the search region for stage two, the tradeoff is acceptable, as it prioritizes anatomical completeness needed for downstream segmentation. These results demonstrate that global-context information, combined with local feature representations, enables ViTs to achieve high segmentation accuracy, as evidenced by the Dice coefficients obtained. Additionally, segmentation performance on this layer yields a PIR of 100%, ensuring that the plexus segmentation is applied within the anatomical region of interest for later accurate segmentation of ganglion cells.
The myenteric plexus region is embedded within the inner layer of the muscularis propria and consists of nerve fibres responsible for several autonomous digestive functions. Unlike the muscularis propria, the myenteric plexus is a smaller, less structured region and thus appears more fragmented across WSIs [26, 27], making segmentation more difficult. The myenteric plexus is of great importance because it is the anatomical site where ganglion cells are present [27], which are needed for future segmentation and identification. The segmentation performance yielded recall and precision scores of 94.8% and 84.2%, respectively. Indicating the model’s robust capability of segmenting most of the pixels that correspond to the plexus region. Additionally, among the tiles identified as plexus, the model produces relatively few false positive segmentations. When oversegmentation does occur, where the predicted plexus region extends slightly beyond the true anatomical boundary, this is generally acceptable because it still maintains a high GIR. In this setting, including a larger area around each plexus reduces the likelihood that a plexus containing ganglion cells will be clipped or partially excluded. This behaviour tends to lower precision, since more pixels are labelled as plexus, but it preserves the downstream objective of ensuring that ganglion-containing regions are retained.
@@
The recall observed (94.8%) for plexus likely reflects instances where entire plexus regions were missed rather than small local errors. This can be examined by recomputing recall after removing completely missed plexus objects from the evaluation, thereby yielding a higher recall value. While some plexus regions are not detected, the very high GIR of 99.7% indicates that the missed regions do not contain ganglion cells and therefore do not influence the success of stage three. It is also possible that the ground truth plexus annotations in these cases were conservative, capturing regions that are anatomically adjacent to the plexus but not functionally relevant. As a result, the model may be omitting less important or marginal areas without affecting the overall goal of preserving ganglion-containing plexus structures. Because the myenteric plexus is embedded within the muscularis propria, producing textural variability [26], this allows ViTs to leverage their inherent self-attention mechanisms, achieving superior results. For instance, the attention layers embedded within the ViT enable the model to distinguish between the thick, continuous, and homogeneous fibres of the muscularis layer and the thin, irregular fibres of the plexus, even when the muscularis propria and the myenteric plexus share the exact colour tones [28]. The ViT computes attention scores between patch pairs, enabling it to identify which patches are similar or different based on the embeddings it produces. This allows the model to distinguish between slight variations in texture. Furthermore, given the fragmented myenteric plexus throughout the WSI, it is discontinuous across several patches, providing ViTs with an opportunity to leverage their inherent understanding of global contextual relationships. Unlike conventional CNNs, which primarily rely on local receptive fields, this may prevent recognizing that spatially separated patches belong to the same structure, leading to fragmented predictions, especially when there are mild structural differences between patches. However, since ViTs operate globally, the model can label patches with slightly different structures as belonging to the plexus region, even when their embeddings are similar. That being said, high segmentation accuracy and high GIR of the myenteric plexus are critical to optimizing ganglion cell segmentation and identification.